### v2

# neutron star; dark matter; gravitational waves; tidal deformability; neutron decay

Abstract
Neutron stars (NSs), among the densest objects in the universe, are exceptional laboratories for investigating the properties of dark matter (DM). Recent theoretical and observational developments have heightened interest in exploring the impact of DM on NS structure, giving rise to the concept of dark matter admixed neutron stars (DANSs). This review examines how NSs can accumulate DM over time, potentially altering their fundamental properties. We explore the leading models describing DM behavior within NSs, focusing on the effects of both bosonic and fermionic candidates on key features such as mass, radius, and tidal deformability. Additionally, we review how DM can modify the cooling and heating processes, trigger the formation of a black hole, and impact gravitational wave (GW) emissions from binary systems. By synthesizing recent research, this work highlights how DANSs might produce observable signatures, offering new opportunities to probe DM’s properties through astrophysical phenomena.
keywords:
neutron star; dark matter; gravitational waves; tidal deformability; neutron decay
[…]
1 Introduction
Compact objects represent a unique class of astrophysical bodies characterized by incredibly high densities and strong gravitational fields. This group includes white dwarfs (WDs), neutron stars (NSs), and black holes (BHs), each formed through the evolutionary endpoints of stellar collapse and distinguished by the degeneracy pressure mechanisms or gravitational forces that prevent further collapse. WDs are sustained by electron degeneracy pressure, limiting their masses to the Chandrasekhar limit of approximately $1.4$ $M_{\mathrm{\odot}}\,$handrasekhar (1939), while NSs are held up by neutron degeneracy pressure and strong nuclear forces Lattimer and Prakash (2004); Baym et al. (2018). BHs, by contrast, lack any such internal support mechanism and represent the ultimate state of stellar evolution, where collapse leads to singularities shielded by event horizons Shapiro and Teukolsky (1983).
Among compact objects, NSs represent ideal environments to probe extreme states of matter, gravitational physics, and, potentially, DM interactions Silk et al. (2010). Primarily composed of densely packed neutrons, NSs have masses ranging from about 1.4 to 2.5 $M_{\mathrm{\odot}}\,$attimer and Prakash (2004); Antoniadis et al. (2013); Özel and Freire (2016), compressed into a radius of roughly 10–12 km Lattimer and Prakash (2007); Steiner et al. (2010); Oertel et al. (2017). The central densities of NSs can exceed several times the nuclear saturation density ($\rho_{\mathrm{nucl}}\sim 2.7\times 10^{14}\,\mathrm{g\,cm^{-3}}$), and can reach values as high as $5\times 10^{15}\,\mathrm{g\,cm^{-3}}$ (Figure 4 in Lattimer and Prakash (2011)) close to the core, orders of magnitude greater than the densities found in ordinary atomic nuclei Lattimer and Prakash (2016); Baym et al. (2018). Such extreme densities lead to intense gravitational fields, with surface gravity
$g_{s}=\frac{M}{R^{2}\sqrt{1-2M/R}}\sim 2\times 10^{12}\,\mathrm{m\;s^{-2}}\,,$ (1)
for an average NS with mass $M=1.4\,M_{\odot}$ and radius $R=10$ km Bejger and Haensel (2004); Haensel et al. (2007) (more than $10^{11}$ times that of Earth). This incredibly strong gravitational field allows for the probing of regimes of physics that are inaccessible to laboratory experiments on Earth.
[…]
The internal structure of NSs is typically modeled in layers. The outermost region is the crust, which extends to a depth of about 1–2 km and consists mainly of nuclei arranged in a crystal lattice immersed in a degenerate electron gas. As one moves deeper, towards the inner crust, neutrons begin to “drip” out of nuclei due to the increasing pressure. The inner core, which makes up most of the star’s mass and volume, is composed of neutron-rich matter, and the exact composition remains an open question. It is theorized that it includes neutrons, protons, electrons, and possibly more exotic particles such as hyperons Logoteta (2021); Oertel et al. (2015); Chatterjee and Vidaña (2016). Furthermore, phase transitions to deconfined quarks Bombaci et al. (2016); Alford et al. (2007); Bombaci and Logoteta (2017) have been hypothesized to occur throughout the whole NS (forming bare quark stars Xu (2003); Page and Usov (2002); Weber et al. (2013)) or near the core (hybrid stars Glendenning (1997); Burgio et al. (2002); Alford and Reddy (2003); Buballa et al. (2004); Alford et al. (2005)). Both ud- and uds-quark matter—commonly referred to as strange quark matter (SQM) Bodmer (1971); Witten (1984); Farhi and Jaffe (1984)—have been proposed Holdom et al. (2018) and studied in NSs Yuan et al. (2022). A recent work by Bai et al. Bai and Chen (2024), however, demonstrates the potential stability of 2-flavor quark matter, while excluding the possibility that (2+1)-flavor quark matter is the ground state in nature.
NS physics can be fully characterized by its Equation of State (EoS) Lattimer and Prakash (2001); Sumiyoshi et al. (2023), describing the relationship between pressure and density. Several models have been developed that are in agreement with observations. However, the exact EoS governing NSs at such high densities remains unknown due to the complex interactions between particles. The EoS is crucial for determining an NS’s mass, radius, and stability against collapse into a BH. Observational constraints on the EoS have been significantly advanced by gravitational wave (GW) detections of binary neutron star (BNS) mergers, like GW170817 Abbott et al. (2018, 2017a, 2017b); Radice et al. (2018) and GW190425 Abbott et al. (2020), which provided direct evidence of the deformability of NSs.
Recently, NSs have been identified as promising candidates for studying DM. Their high densities and temperatures, the strong gravitational field at their surfaces, and their ages (up to $10^{10}\,\mathrm{years}$) allow NSs to collect DM particles over time. This means NSs could serve as natural detectors, helping us learn more about DM, its properties, and its role in the universe Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010); Ciarcelluti and Sandin (2011); Ellis et al. (2018).
Accounting for approximately 27% Aghanim et al. (2020) of the universe’s total density, DM remains a mysterious component that exerts gravitational effects on visible matter, but neither emits nor absorbs light. This makes DM undetectable through traditional electromagnetic observations. First hypothesized in the 1930s by Fritz Zwicky, who noted the unusual motion of galaxies in the Coma Cluster Zwicky (1933), DM’s existence was further supported in the 1970s by observations of flat rotation curves in spiral galaxies Rubin and Ford (1970); Rubin et al. (1978); Mohan and Goswami (2024). These curves suggest that galaxies contain far more mass than what is visible, implying the presence of an unseen gravitational source.
[…]
Experimental efforts to detect DM have been extensive. Astrophysical searches primarily aim to capture signals of GWs and electromagnetic waves from DM interactions in extreme environments, namely, in isolated compact objects and in coalescences of binaries. The experimental techniques vary depending on the mass range under investigation: ultralight DM (<keV) may generate detectable radio or X-ray emissions from NSs, may influence the inspiral phase of GW emission from BNS mergers, and may lead to coherent GW signals through BH superradiance. Light DM (keV–MeV) can be traced through X-ray/gamma-ray emissions from supernovae, temperature anomalies in NSs, or changes in stellar evolution influencing BH populations. Heavy DM (>GeV) may manifest through heating of NSs, DM spikes enhancing annihilation rates near BHs, and merger dynamics detected by future GW observatories. Such methods are described in Baryakhtar et al. (2022), from which we obtain a schematic representation (Figure 1) that summarizes how to probe DM within its wide mass range from astrophysical observations.
Figure 1: Summary of astrophysical searches covering the wide range of DM masses probed by various methods. The schematic representation is adapted from Baryakhtar et al. (2022).
A significant role in the search for DM is played by direct detection experiments, such as those conducted by the XENON1T Aprile et al. (2018, 2019a, 2019b, 2019c, 2021, 2020), DAMA/LIBRA Bernabei et al. (2010), LUX–ZEPLIN Akerib et al. (2014), CRESST Angloher et al. (2016); Abdelhameed et al. (2019), CDMSLite Agnese et al. (2014, 2018), and CYGNUS Vahsen et al. (2020) collaborations, which aim to observe rare interactions between DM particles and atomic nuclei. Meanwhile, indirect detection approaches, such as those conducted by the Fermi Large Area Telescope Atwood et al. (2009), search for signals produced when DM particles potentially annihilate each other. Other promising methods include gravitational lensing studies Clowe et al. (2006); Massey et al. (2007), where the gravitational effect of DM on light from distant galaxies helps map its distribution; and observations of cosmic microwave background (CMB) fluctuations Hooper (2019), which provide indirect evidence of DM’s influence on the early universe’s structure. Works like Battaglieri et al. (2017); Bertone and Hooper (2018) provide comprehensive reviews of the history and current status of DM models and experimental searches.
DM can be accreted into NSs by gravitational capture Weinberg (1978); Marsh (2016); Eilers et al. (2019); Bell et al. (2020); Busoni (2022), which is expected to be more efficient for massive particles such as WIMPs Kouvaris (2008); Feng (2010); Arcadi et al. (2018). Once accumulated, DM could influence the structure, the thermal evolution, and the long-term stability of the new compact object, known as a dark matter admixed neutron star (DANS). Indeed, the presence of DM can trigger different mechanisms. For instance, self-annihilating DM could inject energy into the star, potentially heating it and altering its cooling profile. This effect could be detectable in old NSs, where the heat generated by DM annihilation would cause them to appear hotter than expected Bertoni et al. (2013); Bell et al. (2024). DM interactions can also be constrained by measuring the surface temperature of NSs (usually in systems where no other significant heating mechanism is expected) de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010); Kouvaris (2008); Giannotti et al. (2016): Observations of Cassiopeia A Heinke and Ho (2010), a young NS with a well-documented cooling history, could offer valuable data in this regard Shternin et al. (2011); Giangrandi et al. (2024). Moreover, ultralight DM particles might be emitted from the star, influencing both its spin-down rate and orbital period. These effects can be further constrained by observing GW radiation Hook and Huang (2018); Kumar Poddar et al. (2019, 2020); Seymour and Yagi (2020); Dror et al. (2021); Lambiase and Poddar (2024). Additionally, the possible decay of neutrons into DM particles within the NS could even alter the mass–radius (M–R) relationship Cline and Cornell (2018); Husain and Thomas (2023); Husain et al. (2022).
DM could also leave signatures in the GW signals emitted by BNS mergers. During the inspiral and merger phases of such collisions, the presence of DM could alter the tidal deformability ($\Lambda$) and the gravitational waveform, potentially leading to detectable deviations from the predictions of purely baryonic models Ellis et al. (2018); Kopp et al. (2018); Bhattacharya et al. (2023). This idea gained traction following the detection of GWs from GW170817. Remarkably, the GW190814 Abbott et al. (2020) event involved a $\sim$23232323 $M_{\mathrm{\odot}}\,$ and a compact object with a mass that was challenging to classify: either a high-mass NS or a low-mass BH Most et al. (2020). The uncertainty about the nature of the lighter companion led to speculation that it might be an NS with a significant fraction of DM Das et al. (2021). Future observations by both current detectors and incoming interferometers (such as the Einstein Telescope Hild et al. (2011); Branchesi et al. (2023) and the Cosmic Explorer Abbott et al. (2017); Evans et al. (2023)) could provide crucial insights into the existence of DM within NSs. Finally, ongoing X-ray observations by the NICER mission Gendreau et al. (2012), which aims to measure NS masses and radii with great precision, offer another method for detecting DM effects by looking for any deviations from the mass–radius relationship predicted by purely baryonic frameworks.
So far, the numerous direct and indirect searches for DM have not yet provided any “smoking gun” proof. However, based on the scenarios discussed above, theoretical works have highlighted potential signatures which, if observed, would serve as compelling evidence for the presence of DM in NSs. Bezares et al. Bezares et al. (2019) have identified a characteristic $m=1$ GW mode in the post-merger waveform of DANS mergers, attributed to the one-arm instability Paschalidis et al. (2015) of the remnant. Additionally, they found a faster exponential decay of the $l=|m|=2$ mode compared to BNS coalescences. Despite some degeneracy with ordinary NSs, these features may be distinguishable by interferometers. In a different context, Sun et al. Sun and Wen (2024) have recently proposed a new diagnostic criterion: DM accumulation close to the NS core could lead to DANSs with smaller normal-matter radii but larger tidal deformabilities than ordinary NSs, in which this negative correlation does not arise. As such, if observations reveal two (or more) NSs with very similar mass and one exhibits a smaller radius but a greater $\Lambda$, this would strongly support the existence of DM in NSs. Even though more sophisticated models might modify or exclude this negative correlation, still a significant deviation from the expected M–R and/or $\Lambda$–M relations Ellis et al. (2018) could serve as a valuable clue.
In this work, we present a general overview of DANSs, emphasizing the impact of DM on their properties and examining how these effects align with current experimental bounds. In detail, the review is organized as follows. In Section 2, we describe the accumulation mechanism of DM within NSs, the heating effects from self-annihilating DM (Section 2.1), and the possible formation of BHs that is triggered by the prsence of DM (Section 2.2). Section 3 covers the models that have been studied when describing DANSs. In particular, in Section 3.1 we present the numerical method typically used to solve the stellar structure of DANSs and to compute properties like radius, mass, and tidal deformability; theoretical models accounting for bosonic DM and fermionic DM are presented in Sections 3.2 and 3.3, respectively; while the effects of DM on the mass–radius and tidal–mass relations are explored in Section 3.4. For fermionic DM, we mostly focus on scalar and vector DANSs, namely, DANS where the DM particles interact with each other by exchanging dark scalar or vector fields. The results are then compared to the current observational constraints from ground-based interferometers and NICER. An interesting approach for the computation of the speed of sound inside DANSs is described in Section 3.5, though more detailed models are needed to analyze the influence of DM on the velocity of linear perturbation inside compact objects. Section 4 is dedicated to the potential DM imprint on signals arising from binary NS mergers. The possible decay of SM particles into DM is examined in Section 5. Finally, the main conclusions are summarized in Section 6.
In the following, we consider geometrized units $G=c=1$, unless otherwise stated.
2 Dark Matter Capture and Accumulation
One of the primary mechanisms for DM accumulation within NSs is gravitational capture. As stars orbit around the center of a galaxy, they encounter significant fluxes of DM particles. Interactions between DM and SM particles within stars can lead to energy loss, potentially trapping DM in the star’s gravitational field. This capture process would be particularly efficient due to the extreme density of NSs. Assuming a typical NS of $M\sim 1.5$ $M_{\mathrm{\odot}}\,$and $R\sim 12$ km, the corresponding escape speed at its surface is $v_{\mathrm{esc}}=\sqrt{2M/R}\approx 0.6$, as measured by an observer at rest on the surface of the NS.
The estimation of the cross-section for collision between a DM particle $\chi$ with the NS is explicitly given in Goldman and Nussinov (1989) (the original derivation was aimed at computing the mass flux of WIMPs that may collide with the NS, but an equal approach holds for any DM candidate). For a particle at infinity with velocity $v_{\chi}$ and impact parameter $b$, the closest approach of $\chi$ to NS is $R$ and the scattering cross-section is $\sigma=\pi b^{2}$. Recalling the motion of test particles in the Schwarzschild metric Weinberg (1972), and assuming $v_{\chi}^{2}<<1$ in the derivation of the impact parameter $b=R\frac{v_{\mathrm{esc}}}{v_{\chi}}[1-2M/R]^{-1/2}$, allows us to compute the DM mass capture rate:
$\dot{M}\equiv\frac{dM_{\chi}}{dt}=\rho_{\chi}v_{\chi}\pi b^{2}P_{\mathrm{v}}P_{\sigma}\,,$ (2)
where the overdot denotes the time derivative, $\rho_{\chi}$ is the local DM density, and $P_{\mathrm{v}}$ and $P_{\sigma}$ are the probabilities that a scattered DM particle loses enough energy to be captured and the probability that $\chi$ is scattered, respectively.
The capture rate in Equation (2) depends primarily on the local DM density as well as its velocity, which can vary significantly across different cosmic environments/regions. Close to the Solar System, the local DM density is calculated as $\rho_{\chi}\approx 0.4\,\mathrm{GeV/cm^{3}}\approx 6.7\times 10^{-25}\,\mathrm{g/cm^{3}}$ and the non-relativistic DM velocity, $v_{\chi}\approx 250\,\mathrm{km/s}$ Eilers et al. (2019). Therefore, the resulting DM flux is approximately $\dot{M}_{\mathrm{ss}}\approx 53\,\mathrm{g/s}\approx 3\times 10^{25}\,\,\mathrm{GeV/s}$, as estimated in Goldman and Nussinov (1989).
The probability of DM scattering can be expressed as $P_{\sigma}=1-e^{-\tau}$, where $\tau=\sigma_{\mathrm{n}\chi}/\sigma_{\mathrm{cap}}$, with $\sigma_{\mathrm{n}\chi}$ the nucleon-DM cross-section while $\sigma_{\mathrm{cap}}$ is the capture cross-section. From Bramante and Raj (2024), the latter is
$\sigma_{\mathrm{cap}}=\begin{cases}\sigma_{0}\left(\frac{m_{\mathrm{n}}}{m_{\chi}}\right),&\mathrm{if\;}m_{\mathrm{evap}}<m_{\chi}<m_{\mathrm{n}}\\ \sigma_{0},&\mathrm{if\;}m_{\mathrm{n}}\leq m_{\chi}\leq\mathrm{PeV}\\ \sigma_{0}\left(\frac{m_{\chi}}{\mathrm{PeV}}\right),&\mathrm{if\;}m_{\chi}>\mathrm{PeV}\,,\end{cases}$ (3)
where in Equation (3) $m_{\mathrm{n}}$ is the neutron mass, $\sigma_{0}=\pi\left(\frac{m_{\mathrm{n}}}{M}\right)R^{2}$ the NS geometric cross-section accounting for the effective area in which DM can be captured, and $m_{\mathrm{evap}}\approx 20\frac{T}{10^{3}\mathrm{K}}$ eV ($T$ is the temperature inside the NS) represents a threshold mass below which DM is ineffectively captured. In other words, if $m_{\chi}<m_{\mathrm{evap}}$, the NS’s thermal energy would cause the captured DM to gain enough kinetic energy to escape from the stellar potential well Garani et al. (2019); Garani and Palomares-Ruiz (2022).
[…]
For the sake of completeness, a slightly different derivation of the capture rate of DM particles into compact objects is presented in Bell et al. (2018); Busoni (2022).
Once DM is captured, it undergoes a series of scattering events with SM particles, gradually losing energy until it settles into a stable orbit within the gravitational pull of the star. These scatterings transfer kinetic energy to the NS, contributing to a gradual increase in its temperature. However, this heating effect is not the only factor influencing the star’s thermal dynamics. Rather, the energy influx from DM is counter-balanced by the rate at which the NS emits photons, leading to a delicate equilibrium between heating and cooling. For NSs that are older than approximately one million years, the dominant mechanism for cooling is photon emission. At equilibrium, the two contributions (the heating due to in-falling DM and the cooling coming from photon emission) are equal.
A significant amount of research has explored the possibility of detecting DM by examining the cooling profiles of old NSs ($\tau\gtrsim 10^{6}$$10^{7}$ years) and their surface temperatures. The time evolution of the red-shifted temperature of an NS is given by Hamaguchi et al. (2019)
$C\frac{dT^{\infty}}{dt}=-L_{\nu}^{\infty}-L_{\gamma}^{\infty}+L_{\mathrm{H}}^{\infty}\,,$ (4)
where $C$ is the total heat capacity of the NS; $L_{\nu}^{\infty}$ and $L_{\gamma}^{\infty}$ are the red-shifted luminosities of the neutrino and photon emissions, respectively; and $L_{\mathrm{H}}^{\infty}$ accounts for any mechanisms that contribute to raising the NS’s temperature. In general, heat can be injected into the NS not only through the scattering and annihilation of DM particles, but also via internal processes such as conversion of magnetic, rotational, and chemical energies Gonzalez and Reisenegger (2010). For clarity, it is useful to distinguish these internal mechanisms from the contribution due to DM interactions, expressing $L_{\mathrm{H}}^{\infty}=L_{\mathrm{int}}^{\infty}+L_{\mathrm{DM}}^{\infty}$.
In Equation (4), redshift correction occurs because the NS thermal relaxation is considered complete in $t\lesssim 10^{2}$ years. Beyond this, the red-shifted temperature, defined as $T^{\infty}=T(r)\left[1-2M/R\right]^{1/2}$, becomes constant in the core. Note that the most general definition $T^{\infty}=T(r)e^{\phi(r)}$ (with $e^{2\phi(r)}=g_{tt}(r)$ being the time component of the metric at position $r$) reduces to $T^{\infty}=T(r)\,\left[1-2M/R\right]^{1/2}$ for a spherically symmetric spacetime described by the Schwarzschild metric, as we have implicitly assumed. The neutrino emission is enhanced by Urca processes Gusakov (2002); Page et al. (2004) as well as pair-breaking and -formation processes Flowers et al. (1976); Yakovlev et al. (1999); Kaminker et al. (1999). During the first $t\lesssim 10^{5}$ years following their formation in supernova explosions, NSs primarily cool through neutrino emission from their interior Yakovlev et al. (2001). Hence, the neutrino luminosity is highly suppressed in older NSs. In contrast, the photon luminosity is expected to dominate at $t\gtrsim 10^{5}$ years and is intrinsically linked to the surface temperature $T_{\mathrm{s}}$ by
$L=g\dot{M}=4\pi\sigma R^{2}T_{\mathrm{s}}^{4}\,,$ (5)
where $\sigma=\frac{\pi^{2}k_{\mathrm{B}}^{4}}{60\hbar^{3}c^{2}}=5.670400\times 10^{-8}\;\mathrm{W\,m^{-2}\,K^{-4}}$ is the Stefan–Boltzmann constant (with $k_{\mathrm{B}}=1.380\;\mathrm{J\,K^{-1}}$ being the Boltzmann constant) and $g=GM/R$ accounts for gravitational effects and plays the role of an efficiency factor representing the fraction of the rest-mass energy converted into radiation during the accretion process  Ikhsanov (2003). The second equality, in which $\dot{M}$ is the rate at which mass is added to the star (Equation (2)), holds by assuming that the NS behaves as a black-body radiator, whose luminosity $L=A\sigma T_{\mathrm{s}}^{4}$ follows the Stefan–Boltzmann law Yin et al. (2011); Carroll and Ostlie (2017) for an approximately spherical NS of surface area $A=4\pi R^{2}$. The maximum photon luminosity $L_{\mathrm{max}}$ is obtained when $P_{\sigma_{\chi}}\approx 1$ and $P_{\sigma_{\mathrm{v}}}\approx 1$. Under these conditions, $L_{\mathrm{max}}$ can be computed close to the Solar System ($\sim$10101010 pc) as $L_{\mathrm{max}}\approx 7\times 10^{24}\,\mathrm{GeV/s}$. Remarkably, for an ordinary NS with $R\sim 10$ km, this value corresponds to a superficial, effective temperature of $\sim$2000200020002000 K that may be detectable by cutting-edge infrared telescopes Chatterjee et al. (2023) such as the James Webb Space Telescope Gardner et al. (2006); Kalirai (2018); Rigby et al. (2023).
To investigate the cooling profile, previous works Yakovlev et al. (1999, 2001); Yakovlev and Pethick (2004) have relied on the “minimal cooling” assumption, meaning that nucleons and charged leptons remain in $\beta$-equilibrium throughout the NS’s evolution, with no internal mechanisms contributing to its heating ($L_{\mathrm{int}}^{\infty}=0$). The resulting $T_{s}$ of isolated, old NSs was estimated to be $<$$1000$ K. However, when DM annihilation is included ($L_{\mathrm{DM}}^{\infty}\neq 0$), the predicted surface temperatures rise to $T_{s}\simeq(2-3)\times 10^{3}$ K. This suggests that the measurements of $T_{s}$ of old NSs could provide critical insights and constraints for DM models.
Recently, the $L_{\mathrm{int}}^{\infty}=0$ assumption has been found to be invalid Yanagi (2019), especially for fast-rotating NSs. Due to the gradual loss of rotational energy, the resulting reduction in centrifugal force causes the NS to slightly contract over time, altering the chemical equilibrium conditions among nucleons and leptons. This change cannot be sustained by weak processes, as they become more and more suppressed after $t\gtrsim 10^{6}$ years.
[…]
2.1 Self-Annihilating Dark Matter
In the case of self-annihilating DM (like WIMPs) into SM states, the accumulation of DM inside the NS can lead to significant heating. Generally, the more DM particles are concentrated in the core, the more efficient the annihilation mechanism, though the dependence on the theoretical model describing such an effect is not trivial Bramante and Raj (2024); Bertoni et al. (2013); Garani et al. (2021). If the DM particles have a non-negligible annihilation cross-section, they will annihilate each other, releasing energy into the star’s core and altering its thermal evolution.
The most general equation that governs the evolution of the number $N_{\chi}(t)$ of DM particles accumulated inside an NS (at time $t$) is as follows Gaisser et al. (1986); Zentner (2009); Chen and Lin (2018):
$\frac{dN_{\chi}(t)}{dt}=C_{\mathrm{n}\chi}+C_{\chi\chi}N_{\chi}(t)-C_{\mathrm{ann}}N_{\chi}(t)^{2}\,,$ (6)
where $C_{\mathrm{n}\chi}$ is the DM capture rate via scattering on nucleons, $C_{\chi\chi}$ accounts for DM self-interactions (the complete expressions of $C_{\mathrm{n}\chi}$ and $C_{\chi\chi}$ are derived in Equation (3.7) in Güver et al. (2014) and in Equation (9) in Zentner (2009), respectively), and $C_{\mathrm{ann}}=\langle\sigma v_{\chi}\rangle/V_{\chi}$ is the annihilation rate coefficient, with $\langle\sigma v_{\chi}\rangle$ representing the thermally averaged DM annihilation cross-section and ${V_{\chi}}$ being the volume of the DM core in which annihilation occurs.
Equation (6) can be used in cases where the DM self-annihilation (last term on the right-hand side) is negligible Del Popolo et al. (2020). Conversely, works as Kouvaris (2008); Kouvaris and Tinyakov (2010); Acevedo et al. (2020) have taken into account the self-annihilation but have ignored the DM self-capture (middle term on the right-hand side). In this latter, common framework, Equation (6) reads Bramante and Raj (2024)
$\frac{dN_{\chi}(t)}{dt}\simeq C_{\mathrm{cap}}-C_{\mathrm{ann}}N_{\chi}(t)^{2}\,,$ (7)
where $C_{\mathrm{cap}}=C_{\mathrm{n}\chi}+C_{\chi\chi}\simeq C_{\mathrm{n}\chi}$ describes the total DM accretion via both processes and may be computed through Equation (2).
The simplified Equation (7) has an analytical solution:
$N_{\chi}(t)=\sqrt{\frac{C_{\mathrm{cap}}V_{\chi}}{\langle\sigma v_{\chi}\rangle}}\,\mathrm{tanh}\left(\frac{t}{\tau_{\mathrm{eq}}}\right)\,,$ (8)
with $\tau_{\mathrm{eq}}=\sqrt{V_{\chi}/(C_{\mathrm{cap}}\langle\sigma v_{\chi}\rangle)}$ being the characteristic time-scale for the equilibrium between annihilation and capture to become settled (i.e., $dN_{\chi}(t)/dt\rightarrow 0$ at $t\geq\tau_{\mathrm{eq}}$). The energy released through annihilation can be computed Kouvaris (2008) as $E=C_{\mathrm{ann}}N_{\chi}^{2}m_{\chi}=\frac{\langle\sigma v_{\chi}\rangle}{V_{\chi}}\mathrm{tanh}^{2}\left(\frac{t}{\tau_{\mathrm{eq}}}\right)m_{\chi}$.
Assuming that the gravitational capture is the main mechanism (i.e., the first term on the right-hand side in Equation (7) dominates), the most general estimation of DM capture Bhattacharya et al. (2023) is
$C=1.4\times 10^{20}\;\mathrm{s^{-1}}$$\times\left(\frac{\rho_{\chi}}{0.4\;\mathrm{GeV\,cm^{-3}}}\right)\left(\frac{10^{5}\;\mathrm{GeV}}{m_{\chi}}\right)\left(\frac{\sigma_{n\chi}}{10^{-45}\;\mathrm{cm^{2}}}\right)\times$
×$\times\left(1-\frac{1-e^{-A^{2}}}{A^{2}}\right)\left(\frac{v_{\mathrm{esc}}}{1.9\times 10^{5}\;\mathrm{km\,s^{-1}}}\right)^{2}\left(\frac{220\,\mathrm{km\,s^{-1}}}{v_{\chi}}\right)^{2}\,,$ (9)
where, in addition to the above definitions, $A^{2}=6m_{\chi}m_{n}v_{\mathrm{esc}}^{2}/v_{\chi}^{2}\left(m_{\chi}-m_{n}\right)^{2}$ accounts for inefficient momentum transfers at large $m_{\chi}$. The capture rate in Equation (2.1) corresponds to a number of accumulated DM particles McDermott et al. (2012) of $N_{\chi}\sim 10^{47}$ under standard conditions (whereas the average nuclear baryon number in an NS is $N_{B}\sim 10^{57}$ Berryman et al. (2022)).
To understand whether equilibrium is reached within the lifetime of an NS, one needs to consider the minimum annihilation cross-section required. Following Garani et al. (2021), it is possible to adopt a partial-wave expansion $\langle\sigma v_{\chi}\rangle=a+bv_{\chi}^{2}$ to express the equilibrium conditions for s-wave and p-wave dominations as
$a$ $>7.4\times 10^{-54}\mathrm{cm^{3}/s}\left(\frac{\mathrm{Gyr}}{\tau}\right)^{2}\left(\frac{C_{\mathrm{max}}}{C_{\chi}}\right)\left(\frac{\mathrm{GeV}}{m_{\chi}}\frac{T}{10^{3}\mathrm{K}}\right)^{\frac{1}{2}},$ (10)
$b$ $>2.9\times 10^{-44}\mathrm{cm^{3}/s}\left(\frac{\mathrm{Gyr}}{\tau}\right)^{2}\left(\frac{C_{\mathrm{max}}}{C_{\chi}}\right)\left(\frac{\mathrm{GeV}}{m_{\chi}}\frac{T}{10^{3}\mathrm{K}}\right)^{\frac{1}{2}},$ (11)
where $\tau$ is the average lifetime of an NS (with mass $\sim$1.51.51.51.5 $M_{\mathrm{\odot}}\,$and radius $\sim$12121212 km) and $C_{\mathrm{max}}$ the maximum capture rate.
An intriguing method to investigate DM self-annihilation in NSs is by focusing on the annihilation products. In particular, if the primary products of DM annihilation are feebly interacting mediators between DM and SM particles, these particles could survive long enough to escape the star before decaying into detectable SM states. These mediators, generated through DM capture by NSs or other celestial bodies, could produce observable fluxes of radiation or neutrinos.
2.2 Black Hole Formation
The capture mechanism described in Section 2 allows the DM fraction to increase more and more further into the NS. If DM could self-annihilate into SM particles upon reaching sufficient densities within the NS (like WIMPs), then the total amount of DM that settles close to the core would stop growing at some point. However, certain DM candidates, such as asymmetric DM do not annihilate due to the assumed asymmetry between the number density of particles and antiparticles. Instead, they continue to accumulate over time, creating a central core of DM that contributes to the gravitational field of the star Bramante and Linden (2014). Asymmetric DM is such that DM particles and their antiparticles have unequal abundances. This disproportion means that asymmetric DM particles, unlike symmetric candidates, do not fully annihilate with their counterparts, allowing them to accumulate in celestial bodies over time. This kind of DM is widely studied in astrophysical contexts because it mirrors the asymmetry observed in the visible universe, which is primarily composed of matter (electrons, protons, and nucleons) and not anti-matter. This would possibly provide an explanation for why DM and ordinary matter exist in comparable proportions within the universe Kaplan et al. (2009); Petraki and Volkas (2013); Zurek (2014).
Once captured, DM particles may undergo one of the following:
  • A two-stage process if the DM–nucleon cross-section is not too large Acevedo and Bramante (2019); Acevedo et al. (2021); Kouvaris and Tinyakov (2011). In this scenario, a first thermalization where DM orbits are larger than the size of the hosting NS is followed by a second thermalization once enough energy is lost and the orbit stabilizes to a final thermal radius within the star.
  • Alternatively, in case of a large $\sigma_{n\chi}$, all captured DM particles will slow down rapidly in the outer shells of the star and only drift, on much longer timescales, towards the core under the influence of a viscous drag force Gould et al. (1990); Starkman et al. (1990); Mack et al. (2007); Bramante et al. (2020).
In both scenarios, the thermal radius, as well as the thermalization time, can be estimated as Goldman and Nussinov (1989); Kouvaris and Tinyakov (2010); Bertoni et al. (2013)
$r_{\mathrm{th}}$ $=2.2\,\mathrm{m}\,\left(\frac{T}{10^{5}}\right)^{\frac{1}{2}}\left(\frac{\mathrm{GeV}}{m_{\chi}}\right)^{\frac{1}{2}}\,,$ (12)
$t_{\mathrm{th}}$ $=0.2\,\mathrm{yr^{-1}}\left(\frac{m_{\chi}}{\mathrm{TeV}}\right)^{2}\left(\frac{\sigma}{10^{-43}\,\mathrm{cm^{2}}}\right)^{-1}\left(\frac{T}{10^{5}}\,\mathrm{K}\right)^{-1},$ (13)
by assuming that the typical DM particle becomes non-relativistic after having been captured and having lost most of its energy.
At some point a critical threshold may be reached, meaning that enough DM has been accumulated into the thermalized region within the NS to trigger the collapse to a BH. This condition can be exploited to limit the DM parameter space as in Fig. 2. Such a process is strongly dependent on the properties of DM particles, especially their mass, interaction cross-section, possible self-interactions, and the degeneracy pressure they experience. Assuming the presence of a WIMP sphere, Kouvaris et al. Kouvaris and Tinyakov (2011) estimated that the onset of the gravitational collapse occurs when the total number $N$ of accumulated WIMPs is $N\sim 5\times 10^{43}\,\left(m_{\chi}/\mathrm{TeV}\right)^{-3}$. For such a value, the potential energy of (semi-relativistic) WIMPs exceeds the Fermi momentum, and therefore Pauli blocking cannot prevent the collapse anymore. The properties of the DM candidate may also influence the likelihood and/or the speed of the gravitational collapse. For example, if the captured DM consists of bosons that form a Bose–Einstein condensate within the star, it would exert minimal pressure and the collapse would be further accelerated Kouvaris and Tinyakov (2011); McDermott et al. (2012); Kouvaris and Tinyakov (2013); Jamison (2013); Garani et al. (2022).
According to Acevedo et al. (2021), in order for DM to trigger the conversion of the NS, the following conditions must hold:
  1. 1.
    The sound crossing time $t_{\mathrm{s}}=r_{\mathrm{th}}/c_{\mathrm{s}}$ (where $c_{\mathrm{s}}$ is the sound speed in a DANS modeled as a perfect fluid) should be greater than the DM freefall time $t_{\mathrm{ff}}$, which is computed by considering the free fall of a test mass onto an object of mass $\frac{4}{3}\pi\rho_{\chi}r_{\mathrm{th}}^{3}$. The condition is
    $\frac{3}{\sqrt{4\pi\rho_{\mathrm{NS}}}}=t_{\mathrm{s}}\geq t_{\mathrm{ff}}=\sqrt{\frac{3\pi}{32\rho_{\chi}}}\,,$ (14)
    which is also known as Jeans stability (see Herrera and Santos (1997); Chavanis (2011) and references therein for similar analysis in other contexts) and is approximately realized when the density of the DM sphere is greater than the solar or terrestrial core density Acevedo et al. (2021).
  2. 2.
    The mass of captured DM should overcome a threshold mass, given as
    $M_{\mathrm{coll}}=\sqrt{\frac{3T_{\mathrm{NS}}^{3}}{\pi m_{\chi}^{3}\rho_{\mathrm{NS}}}}\approx 2.2\times 10^{54}\,\mathrm{GeV}\,\left(\frac{10^{7}\mathrm{GeV}}{m_{\chi}}\right)^{\frac{3}{2}}\,,$ (15)
[…]
    For $M_{\mathrm{DM}}<M_{\mathrm{coll}}$, namely, prior to collapse, the DM particles are in thermal equilibrium within a dense sphere of radius $r_{\mathrm{th}}$ at temperature $T_{\mathrm{NS}}$ of the hosting NS’s core ($r_{\mathrm{th}}$ and $T_{\mathrm{NS}}$ are related through the virial theorem $2\langle K\rangle=-\langle U\rangle$, where in turn, $\langle K\rangle=3/2\cdot T_{\mathrm{th}}$ and $\langle U\rangle=\frac{4}{3}\pi r_{\mathrm{th}}^{2}\rho_{\mathrm{NS}}m_{\chi}$). Once the condition $M_{\mathrm{DM}}>M_{\mathrm{coll}}$ is reached, the DM core can no longer be stabilized and the collapse to a BH is irreversibly triggered.
    In addition, the central DM sphere must be sufficiently massive such that any additional contributions from self-interactions or degeneracy pressure are not enough to stabilize it and prevent collapse. More precisely, the critical mass beyond which collapse occurs should be defined as $M_{\mathrm{cr}}=\mathrm{max}[M_{\mathrm{coll}},\,M_{i}]$, where $M_{i}$ Bramante and Raj (2024) represents the maximum mass that can be stabilized by interactions and is given by
    $M_{i}\equiv\begin{cases}M_{f}\sim M_{\mathrm{pl}}^{3}/m_{\chi}^{2}\sim\,M_{\odot}\left(\frac{\mathrm{GeV^{2}}}{m_{\chi}^{2}}\right)\,\;\;&\mathrm{for\;\;fermionic\;\;DM}\,,\\ M_{b}\sim\frac{2M_{\mathrm{pl}}^{2}}{m_{\chi}}\left(1+\frac{\lambda}{32\pi}\frac{M_{\mathrm{pl}}^{2}}{m_{\chi}^{2}}\right)^{1/2}\;\;&\mathrm{for\;\;bosonic\;\;DM}\,,\end{cases}$ (16)
    with $M_{\mathrm{pl}}=1.2209\times 10^{19}$ GeV being the Planck mass and $\lambda$ the coupling constant of the possible boson repulsive self-interaction ($\mathcal{L}\supset-(\lambda/4!)\chi^{4}$).
Once the BH is formed close to the center, its time evolution depends on competing effects coming from accretion (of both nuclear and DM) and Hawking radiation. Specifically, a small enough BH could even evaporate completely; in contrast, a large BH is likely to accumulate surrounding material and possibly convert the whole NS. The rate at which the mass of the BH, $M_{\mathrm{BH}}$, changes has been derived in Garani et al. (2019); Acevedo et al. (2021):
$\small\frac{dM_{\mathrm{BH}}}{dt}=\dot{M}^{\mathrm{(NS\;acc)}}+\dot{M}^{(\mathrm{DM\;acc)}}-\dot{M}^{(\mathrm{Haw\;rad})}=\frac{4\pi\rho_{\mathrm{NS}}M_{\mathrm{BH}}^{2}}{c_{s}^{3}}+e_{\chi}m_{\chi}C_{\mathrm{cap}}-\frac{f(M_{\mathrm{BH}})}{M_{\mathrm{BH}}^{2}}\,,$ (17)
where $e_{\chi}\in[0,1]$ is an efficiency factor accounting for additional DM falling into the newborn BH, $C_{\mathrm{cap}}$ is the DM capture rate defined in Equation (7), and $f(M_{\mathrm{BH}})$ is the Page factor Page (1976) that relates the SM degrees of freedom radiated from the BH to the gray-body corrections MacGibbon and Webber (1990); MacGibbon (1991).
[…]
Overall, if $\frac{dM_{\mathrm{BH}}}{dt}>0$, the BH grows in the NS; otherwise, $\frac{dM_{\mathrm{BH}}}{dt}<0$ yields a partial or total evaporation. Remarkably, the BH’s fate is almost entirely determined by its initial mass McDermott et al. (2012); Kouvaris and Tinyakov (2011). From Equation (17), Garani et al. Garani et al. (2019) obtained lower bounds for the mass of different DM candidates that would cause the evaporation of such a BH as
$m_{\chi}\gtrsim\begin{cases}3\times 10^{6}\mathrm{\;GeV\;for\;bosons\;that\;do\;not\;form\;a\;BEC;}\\ 16\mathrm{\;GeV\;for\;bosons\;that\;do\;form\;a\;BEC;}\\ 10^{10}\mathrm{\;GeV\;for\;fermions.}\end{cases}$ (18)
Experimental searches for DM signatures in NSs have increasingly focused on identifying anomalies in NS populations and characteristics, particularly in regions of high DM density, such as the galactic center or within dwarf spheroidal galaxies. Several experimental areas are worth focusing on to search for BH formation due to DM accretion in NSs. The existence of old pulsars (in our galaxy, many pulsars with ages on the order of several billion years have been observed Smith et al. (2023); most of these are located within 1 kpc, where the DM density is approximately $\rho_{\chi}=0.3\,\mathrm{GeV/cm^{3}}$) in regions with known DM density can serve as a constraint on the amount of DM captured by NSs Kouvaris (2008); Kouvaris and Tinyakov (2010); Bramante and Elahi (2015) (see left panel in Figure 2). These considerations about old NSs also significantly limit the DM parameter space and the DM–nucleon cross-section (right panel in Figure 2). These exclusion curves have been discussed in McDermott et al. (2012); Bertoni et al. (2013); Garani et al. (2019), who also provided a similar analysis for muons. Additionally, the observed absence of millisecond pulsars in the Galactic Center of the Milky Way Perna et al. (2003); Dexter and O’Leary (2014); Suresh et al. (2022) has been hypothesized to result from NS implosions triggered by DM accumulation de Lavallaz and Fairbairn (2010); Fuller and Ott (2015); Bramante (2015). A possible source of r-process elements may be supernovae that follow NSs’ collapse Bramante and Linden (2016); Bramante et al. (2018).
3 Dark Matter Admixed Neutron Star Models
DANSs are astrophysical environments where a fraction of DM has been accreted into the stellar volume through the physical mechanisms described in Section 2. To model the influence of DM on NSs, one must first consider how the two types of matter—the baryonic and the dark sectors—interact. While a number of interactions are possible, including direct coupling through forces beyond gravity, null experimental evidence has placed strict limits on non-gravitational interactions between these components. In particular, DM direct detection experiments and the Bullet Cluster Clowe et al. (2006); Randall et al. (2008); Marrodán Undagoitia and Rauch (2016) have constrained the DM-BM cross-section to be many orders of magnitude lower than the typical nuclear one $\sigma_{\mathrm{DM-BM}}\sim 10^{-45}\,\mathrm{cm^{2}}<<\sigma_{\mathrm{nucl}}\sim 10^{-24}\,\mathrm{cm^{2}}$. Figure 3 shows the latest experimental limits on the nucleon cross-section for spin-independent DM–nucleon interactions (top panel) and spin-dependent interactions (bottom panel) derived from the publicly available data of the LUX–ZEPLIN collaboration Aalbers et al. (2024). Interestingly, Kouvaris et al. Kouvaris and Tinyakov (2011) pointed out that for massive DM candidates, the spin-dependent cross-section can be further constrained from the existence of NSs in globular clusters (see Figure 2 in Kouvaris and Tinyakov (2011), while Figure 1 represents a similar analysis when WDs are assumed to capture DM), that is, by looking at WIMPs’ accretion into NSs or in a progenitor supramassive star that later collapses to an NS.
Figure 3: Upper limits (68% and 95% CL) on the DM–nucleon cross-section for spin-independent (SI) interactions (top panel) as well as spin-dependent (SD) DM–neutron (bottom left) and DM–proton (bottom right) interactions from the latest public data release of the LUX–ZEPLIN collaboration Aalbers et al. (2024) (from which the plots are adapted). The results are obtained by combining the first 60-live-day exposure (WS2022) with a new 220-live-day exposure campaign (WS2024). Experimental limits derived from WS2022 only Aalbers et al. (2023), LUX Akerib et al. (2017), Xenon1T Aprile et al. (2018), XENONnT Aprile et al. (2023), DEAP3600 Ajaj et al. (2019), and PICO–60 Amole et al. (2019) are also shown.
These experimental constraints significantly reduce the likelihood of strong coupling, guiding theoretical efforts toward more conservative approaches. In other words, from the perspective of astrophysical dense environments as NSs, this generally means that any interaction between DM and the ordinary matter of the NS is negligibly small Nelson et al. (2019); Deliyergiyev et al. (2019). As a consequence, the most common and physically motivated framework to solve DANSs’ structure is the two-fluid model; namely, a two-component version of the Tolman–Oppenheimer–Volkoff (TOV) equations Oppenheimer and Volkoff (1939); Tolman (1939), describing the hydrodynamical equilibrium of a spherically symmetric, non-rotating star under general relativity (GR). In DANS models, the two components are the baryonic (neutron-rich) matter and the DM. They are treated as separate fluids, interacting only gravitationally. The two-fluid approach is particularly advantageous for several reasons. First, it respects the minimal interaction scenario, supported by current cosmological and astrophysical observations as well as direct detection experiments. Second, this model is computationally accessible, allowing for simulations of NS structures while incorporating distinct equations of state for both baryonic matter and DM that separately maintain their own thermodynamic properties (i.e., pressure and energy density). Additionally, the two-fluid model enables the exploration of a wide variety of DM candidates (Sections 3.2 and 3.3 below). Each of these modifies the stellar structure differently, but the two-fluid framework can be adapted to accommodate these different scenarios.
Overall, the theoretical models analyzing the influence of DM on NS properties must account for both the gravitational coupling between the two components as well as any potential self-interactions within the dark sector. Research in this area has employed different approaches to describe either the baryonic or the dark fluid. While variations in the EoS for BM yield only minor differences due to the stringent observational constraints they must satisfy (see Burgio et al. (2021) for a comprehensive review on nuclear EoSs), the uncertain nature of DM leads to the consideration of a variety of scenarios, in which DM is either gravitationally captured by the NS or is embedded from the time of the star’s formation. Regarding the nature of DM, it can consist of either bosons Chavanis and Harko (2012); Giangrandi et al. (2023); Rafiei Karkevandi et al. (2021); Karkevandi et al. (2022) or fermions Narain et al. (2006); Goldman et al. (2013). In the former case, the DM candidates form a Bose–Einstein condensate close to the star’s center; in the latter, particles may be free or interact through the exchange of scalar and vector dark mediators. Detailed models include work by Leung et al. Leung et al. (2011) and Ivanytskyi et al. Ivanytskyi et al. (2020), who modeled DM as a free Fermi gas, while Ellis et al. Ellis et al. (2018) explored both self-interacting bosonic DM, which can form a Bose–Einstein condensate, and asymmetric fermionic DM. Further studies by Karkevandi et al. Karkevandi et al. (2024); Shakeri and Karkevandi (2024) and Liu et al. Liu et al. (2023, 2024) have examined the effects of both bosonic and fermionic DM. Additional research has expanded the range of DM models: Jockel et al. Jockel and Sagunski (2024) studied fermionic Proca stars; Diedrichs et al. Diedrichs et al. (2023) focused on the properties of ultralight DM; mirror DM was investigated in Ciancarella et al. (2021); Berezhiani et al. (2021); Kain (2021); Berezhiani (2022); strange stars with mirror DM have been studied in Yang et al. (2021); Yang and Pi (2024); Yang et al. (2024); finally, Rezaei Rezaei (2023) analyzed fuzzy DM. For the sake of completeness, we report that alternative ideas have been proposed and investigated by Shahrbaf et al. Shahrbaf et al. (2022, 2024), who assumed (bosonic) DM consists of stable sexa-quark states; scalarization of DM inside NSs within modified theories of gravity has been considered in Chen et al. (2015); Morisaki and Suyama (2017); Sotani and Kokkotas (2017); Davoudiasl (2024); and superfluid models have been applied to either the DM component Reddy and Zhou (2022) or the nuclear matter Acevedo et al. (2020); Fujiwara et al. (2024).
It is also worth mentioning that recently some authors made use of a single-fluid framework. Works like Panotopoulos and Lopes (2017); Das et al. (2019, 2021a, 2021b, 2021); Sen and Guha (2021); Lenzi et al. (2023); Dutra et al. (2022); Flores et al. (2024); Hong and Ren (2024); Thakur et al. (2024) do not consider gravitational interactions, rather they exploit a Higgs portal mechanism to couple DM and BM. DANSs with feeble interactions between DM and hadronic matter have been studied in Guha and Sen (2021, 2024); Sen and Guha (2021). Alternatively, Refs. Motta et al. (2018); Husain et al. (2022); Shirke et al. (2023, 2024); Bastero-Gil et al. (2024) assumed that DANSs may form from neutron decay into dark matter (see Section 5). These several different approaches highlight the extensive exploration of DM’s potential impacts on NSs’ characteristics.
3.1 Formalism
The general relativistic metric describing a static and spherically symmetric spacetime is
$ds^{2}=-e^{\nu(r)}dt^{2}+\frac{dr^{2}}{1-2m(r)/r}+r^{2}(d\theta^{2}+sin^{2}\theta d\phi^{2})\,,$ (19)
where $m(r)$ is the enclosed mass within radius $r$ and $\nu(r)$ is the metric function that decouples for a static metric. The baryonic and dark sectors are usually modeled as perfect fluids, whose stress–energy tensors are
$T_{i}^{\mu\nu}=(\epsilon_{i}+P_{i})u^{\mu}u^{\nu}+P_{i}\eta^{\mu\nu}\,,$ (20)
where $\epsilon_{i}$ and $P_{i}$ ($i=\mathrm{BM},\mathrm{DM}$) are the energy densities and the pressures of the two fluids separately. From such definitions, the total stress–energy tensor of the DANS is computed as their sum $T_{\mathrm{tot}}^{\mu\nu}=T_{\mathrm{BM}}^{\mu\nu}+T_{\mathrm{DM}}^{\mu\nu}$. Since the two fluids only interact gravitationally, the conservation law holds separately for each fluid, i.e., $\nabla_{\mu}T_{i}^{\mu\nu}=0$. Then, one can obtain the coupled TOV equations by solving the Einstein field equations for the metric in Equation (19). They describe the hydrostatic equilibrium of the DANS and they take the form
$\frac{dP_{i}}{dr}$ $=-(P_{i}+\epsilon_{i})\frac{4\pi r^{3}P_{\mathrm{tot}}+m(r)}{r(r-2m(r))}\,,$ (21)
$\frac{dm(r)}{dr}$ $=4\pi\epsilon_{\mathrm{tot}}r^{2}\,,$ (22)
where $P_{\mathrm{tot}}=P_{\mathrm{BM}}+P_{\mathrm{DM}}$, $\epsilon_{\mathrm{tot}}=\epsilon_{\mathrm{BM}}+\epsilon_{\mathrm{DM}}$ are the total pressure and energy density, respectively.
Equations (21) and (22) form a system of equations that is closed by the equation of state $P_{i}\equiv P_{i}(\epsilon_{i})$, usually expressed in the barotropic approximation Bauswein et al. (2012). To numerically solve such a system, one first needs to specify the central density $(\epsilon^{\mathrm{c}})$ of the star that acts as a boundary condition. Equations (21) and (22) are then integrated outward, starting from the core of the star (where $r=0$) and reaching the surface (defined as the locus of points where $P_{\mathrm{tot}}(r=R)=0$; this condition implicitly defines the star’s physical radius and ensures that the solution is consistent with the hydrostatic equilibrium of the DANS).
In addition, it is important to note that solving the equations for two-fluid systems requires specifying the fraction $f_{\mathrm{DM}}$. While a few studies Das et al. (2022) express it in terms of the central densities of the two stars, i.e., $f_{\mathrm{DM}}=\epsilon_{\mathrm{DM}}^{\mathrm{c}}/\epsilon_{\mathrm{BM}}^{\mathrm{c}}$, most of the works Karkevandi et al. (2022); Panotopoulos and Lopes (2017); Ivanytskyi et al. (2020) define it as the ratio between the DM mass and the total mass of the star, i.e., $f_{\mathrm{DM}}=M_{\mathrm{DM}}/M_{\mathrm{star}}=M_{\mathrm{DM}}/(M_{\mathrm{BM}}+M_{\mathrm{DM}})$. Fixing this fraction allows for the computation of a one-dimensional mass–radius relation, facilitating comparisons with the M–R curves of ordinary NSs and validating or refuting models. However, this approach can be significantly model-dependent, as it assumes a specific mechanism for DM accumulation. The capture details are sensitive to the nature of DM self-interactions, interactions between DM and standard particles, and other uncertain astrophysical factors. A more consistent approach would treat the DM fraction as a variable parameter, leading to a two-dimensional space of solutions rather than a single curve Hippert et al. (2023). While this enhances the exploration of the stellar properties impacted by the DM fraction, it complicates direct comparisons with NS models. This complexity is one reason many authors choose to fix $f_{\mathrm{DM}}$ in their analyses.
Another key observable that has emerged from the study of NSs, especially in the GW era, is the tidal deformability, a measure of how easily an NS’s shape is distorted by an external gravitational field. This property is particularly important in the context of BNS mergers, where tidal interactions between the stars leave an imprint on the GW signal Radice et al. (2018); Raithel et al. (2018); Zhao and Lattimer (2018). The dimensionless tidal deformability parameter Hinderer (2008); Ellis et al. (2018) is
$\Lambda=\frac{2}{3}k_{2}\left(\frac{M}{R}\right)^{-5}\,,$ (23)
where $k_{2}$ is the $\ell=2$ tidal Love number Flanagan and Hinderer (2008); Damour and Nagar (2009), with a typical range from $0.05$ to $0.15$ Hinderer et al. (2010); Postnikov et al. (2010). From Hinderer (2008), the Love number is computed as
$k_{2}$ $=\frac{8C^{5}}{5}(1-2C)^{2}[2+2C(y_{R}-1)-y_{R}]\Big{\{}2C(6-3y_{R}+3C(5y_{R}-8))+$ (24)
$4C^{3}\big{[}13-11y_{R}+C(3y_{R}-2)+2C^{2}(1+y_{R})\big{]}\times$
×$\times 3(1-2C)^{2}\big{[}2-y_{R}+2C(y_{R}-1)\big{]}\mathrm{\ln}(1-2C)\Big{\}}^{-1}\,,$
where $C=M/R$ is the compactness parameter and $y_{R}\equiv y(r=R)$ is the solution of the two-fluid differential equation
$r\frac{dy(r)}{dr}+y(r)^{2}+y(r)F(r)+r^{2}Q(r)=0\,,$ (25)
where in turn,
$F(r)$ $=\frac{r-4\pi r^{3}(\rho_{tot}-P_{tot})}{r-2m(r)}\,,$ (26)
$Q(r)$ $=\frac{4\pi r[5\rho_{tot}+9P_{tot}+\sum_{i}\frac{\rho_{i}+P_{i}}{\partial P_{i}/\partial\rho_{i}}-\frac{6}{4\pi r^{2}}}{r-2m(r)}-4\left[\frac{m(r)+4\pi r^{3}P_{tot}}{r^{2}(\-2m(r)/r)}\right]^{2}\,.$ (27)
3.2 Bosonic DM
To model bosonic DM, a complex scalar field with mass $m_{\chi}$ and chemical potential $\mu_{\chi}$ is considered. These particles are not subject to the Pauli exclusion principle and they can accumulate in close proximity. At sufficiently low temperatures, bosonic DM exists in the form of a Bose–Einstein condensate (BEC) Li et al. (2012); Arbey et al. (2003); Suárez et al. (2014). In the absence of interactions, a BEC has zero pressure and is mechanically unstable against gravitational compression. To stabilize the BEC, a repulsive interaction mediated by a real vector field $\omega_{\mu}$ (that couples with $\chi$) must be introduced. The minimal Lagrangian accounting for this interaction has been discussed in detail by Giangrandi et al. Giangrandi et al. (2023). This model is equivalent to a massive $U(1)$ gauge theory (like scalar electrodynamics with a massive photon). Thus, the Lagrangian includes the mass and kinetic terms of the fields $\chi$ and $\omega_{\mu}$ that are coupled through the covariant derivative $D_{\mu}=\partial_{\mu}-ig\omega_{\mu}$ ($g$ is the coupling constant):
$\mathcal{L}_{\mathrm{B}}=(D_{\mu}\chi)^{*}D^{\mu}\chi-m_{\chi}^{2}\chi^{*}\chi-\frac{\Omega_{\mu\nu}\Omega^{\mu\nu}}{4}+\frac{m_{\omega}^{2}\omega_{\mu}\omega^{\mu}}{2}\,.$ (28)
In Equation (28), $m_{\chi}$ and $m_{\omega}$ are the masses of the scalar and the vector field, respectively, and $\Omega_{\mu\nu}=\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}$ is the vector field tensor.
It is worth noting that other studies have examined variations to this model. Nelson et al. Nelson et al. (2019) and Rutherford et al. Rutherford et al. (2023) investigated asymmetric bosonic DM, considering not only repulsive self-interactions but also an additional term $g_{\mathrm{B}}\omega_{\mu}J^{\mu}_{\mathrm{B}}$, where $g_{\mathrm{B}}$ represents the interaction strength of $\omega_{\mu}$ with the SM baryon number current $J^{\mu}_{\mathrm{B}}$ and the whole term accounts for small attractive interactions. Alternatively, Karkevandi et al. Karkevandi et al. (2022) considered a self-interaction potential $V(\chi)=\lambda|\phi|^{4}$, where $\lambda$ is the dimensionless coupling constant of the scalar field forming the BEC. This modification leads to a different EoS of the form
$P=\frac{m_{\chi}^{4}}{9\lambda}\left(\sqrt{1+\frac{3\lambda}{m_{\chi}^{4}}}-1\right)\,.$ (29)
Overall, while the qualitative impact of bosonic DM on DANSs remains similar, different models may be adopted due to the unknown nature of DM.
The Lagrangian in Equation (28) gives rise a Noether current that corresponds to the invariance of the action under global $U(1)$ transformations $j^{\mu}=i(\chi^{*}\partial\chi-\chi\partial\chi^{*})+2g\omega^{\mu}\chi^{*}\chi$. Note that the last term in the right-hand side does not contribute to the EoS. Indeed, for the computation of $\epsilon$ and $\rho$, the relevant quantity is the density of conserved charge associated with the DM particles. This density is obtained by averaging the zeroth component of $j^{\mu}$. When performing this average, $2g\omega^{\mu}\chi^{*}\chi$ provides a vanishing contribution.
By assuming cold NSs ($T\sim 0$), the bosonic field totally converts into a BEC, and thermal fluctuations are suppressed. This allows for the use of a relativistic mean-field (RMF) approach Shen et al. (1998); Diener (2008) to derive the EoS. In this framework, the vector field $\omega_{\mu}$ can be replaced by its constant mean value $\langle\omega^{\mu}\rangle$. After some algebra, the expressions of pressure and energy density are obtained from the Euler–Lagrange equations as
$P_{\chi}=\frac{m_{\mathrm{I}}^{4}}{4}\left(m_{\chi}^{2}-\mu_{\chi}\sqrt{2m_{\chi}^{2}-\mu_{\chi}^{2}}\right)\,,$ (30)
$\epsilon_{\chi}=\frac{m_{\mathrm{I}}^{4}}{4}\left(\frac{\mu_{\chi}^{3}}{\sqrt{2m_{\chi}^{2}-\mu_{\chi}^{2}}}-m_{\chi}^{2}\right)\,,$ (31)
with $m_{\mathrm{I}}$ being a fictional mass that controls the strength of the coupling ($m_{\mathrm{I}}\rightarrow\infty$ when $g\rightarrow 0$ and vice-versa). It can be shown Giangrandi et al. (2023) that no values of $m_{\mathrm{I}}$ lead to infinite density, but, to remain consistent with experimental observations from the Bullet Cluster, a conservative bound of $m_{\mathrm{I}}\,\mathrm{[MeV]}>18.24\sqrt[4]{m_{\chi}\mathrm{[MeV]}}$ is needed to keep the parameter space as broad as possible.
3.3 Fermionic DM
Fermionic DM, that must obey the Pauli exclusion principle, is described by a Dirac field $\Psi$. Some studies have modeled DM as a relativistic Fermi gas of non-interacting particles with mass $m_{\chi}$. Starting from the Lagrangian density of a Dirac field, the corresponding EoS for a cold DANS reads Shapiro and Teukolsky (1983); Glendenning (1997)
$\epsilon$ $=\frac{m_{\Psi}^{4}}{2}\,\eta(x)=\,\frac{m_{\Psi}^{4}}{2}\frac{1}{8\pi^{2}}\left[x\sqrt{1+x^{2}}(1+2x^{2})-\mathrm{ln}(x+\sqrt{1+x^{2}})\right]\,,$ (32)
$P$ $=\frac{m_{\Psi}^{4}}{2}\,\Xi(x)=\frac{m_{\Psi}^{4}}{2}\,\frac{1}{8\pi^{2}}\left[x\sqrt{1+x^{2}}\left(\frac{2}{3}x^{2}-1\right)+\mathrm{ln}(x+\sqrt{1+x^{2}})\right]\,,$ (33)
where $x=k_{\mathrm{F}}/m_{\Psi}$ is the dimensionless Fermi momentum.
[…]
$\mathcal{L}_{\mathrm{F}}$ $=\bar{\Psi}(x)\left\{\gamma_{\mu}\left[i\partial^{\mu}-g_{v}V^{\mu}(x)\right]-\left[m_{\Psi}-g_{s}\phi(x)\right]\right\}\Psi(x)+$
$+\frac{1}{2}\big{[}\partial_{\mu}\phi(x)\partial^{\mu}\phi(x)-m_{s}^{2}\phi^{2}(x)\big{]}-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{v}^{2}V_{\mu}(x)V^{\mu}(x),$ (34)
where $g_{v}$ and $m_{v}$ are, respectively, the coupling constants and the mass of the vector field $V_{\mu}$, $g_{s}$ and $m_{s}$ those of the scalar field $\phi$ and $V_{\mu\nu}=\partial_{\mu}V_{\nu}(x)-\partial_{\nu}V_{\mu}(x)$.
The exchanges of mesons described by the two terms $g_{s}\phi\bar{\Psi}\Psi$ and $g_{v}\bar{\Psi}\gamma_{\mu}\Psi V^{\mu}$ give rise to an effective Yukawa potential,
$V(r)=\frac{g_{v}^{2}}{4\pi}\frac{e^{-m_{v}r}}{r}-\frac{g_{s}^{2}}{4\pi}\frac{e^{-m_{s}r}}{r}\,,$ (35)
whose behavior is governed by the couplings $g_{i}$ and the masses $m_{i}$ ($i=s,v$) of the mediators. There is no unique choice for these values, as the DM Lagrangian leads to an EoS that depends only on the ratios $c_{i}=g_{i}/m_{i}$. Typically $c_{i}\sim\mathcal{O}(1-15)\,\mathrm{GeV^{-1}}$, with $c_{s}<c_{v}$ so that the potential is attractive at large distances and repulsive at short distances Xiang et al. (2014); Das et al. (2022).
Within the RMF approximation, one can replace the mediator field operators with their ground state expectation values $\left\langle\phi(x)\right\rangle=\phi_{0}$, $\left\langle V_{\mu}(x)\right\rangle=V_{0}$. These values are computed from the Euler–Lagrange equations and depend on the effective mass of the DM fermion candidate, which reduces due to the scalar coupling as $m_{\Psi}^{*}=m_{\Psi}-g_{s}\phi_{0}$.
Finally, the energy density and the pressure corresponding to the Lagrangian in Equation (34) are given by
$\epsilon$ $=\frac{{m_{\Psi}^{*}}^{4}}{\pi^{2}}\eta(x)+\frac{g_{v}^{2}}{2m_{v}^{2}}\rho_{\Psi}^{2}+\frac{g_{s}^{2}}{2m_{s}^{2}}(m_{\Psi}-m_{\Psi}^{*})^{2}\,,$ (36)
$P$ $=\frac{{m_{\Psi}^{*}}^{4}}{3\pi^{2}}\Xi(x)+\frac{g_{v}^{2}}{2m_{v}^{2}}\rho_{\Psi}^{2}-\frac{g_{s}^{2}}{2m_{s}^{2}}(m_{\Psi}-m_{\Psi}^{*})^{2}\,,$ (37)
where the functions $\eta(x)$ and $\Xi(x)$ correspond to the “free” terms defined in Equations (32) and (33). Equations (36) and (37) represent the equation of state for fermionic DM that interacts both with scalar and vector mediators.
3.4 Mass–Radius and Tidal–Mass Relations
In this subsection, we summarize the main effects of DM on the mass–radius and tidal–mass relations of DANSs. Indeed, the existence of a DM component can significantly alter the NS’s properties. This happens because the accreted DM can either form a dense core within the NS or a halo surrounding the compact object Kain (2021). Unless much extended halos are formed, DM accumulation typically reduces a DANS’s mass. In an NS, all the BM contributes to the internal pressure, whereas in a DANS the DM fraction typically has a negligible effect on this support. As a result, an NS composed solely of BM could exert more pressure and sustain a higher mass.
In detail, we compare the M–R (Figure 4) and $\Lambda$–M (Figure 5) diagrams for nuclear equations of state (left panels) with DANS models (right panels). Fermionic, interacting DM is assumed to accumulate in the NSs, as studied in Grippa et al. (2024).
In the right panels of Figure 4, fixed masses of dark particles are considered. In particular a nucleon-like DM candidate (this is particularly interesting because such a DM particle may be related to the mirror particle of ordinary BM Ciarcelluti and Sandin (2011); Khlopov et al. (1991)) with $m_{\Psi}=1$ GeV and ultralight mediators with $m_{s}=10^{-14}$ eV and $m_{v}=10^{-6}$ eV. These choices result in $c_{s}\sim 5\,\mathrm{GeV^{-1}}$ and $c_{v}\sim 10\,\mathrm{GeV^{-1}}$, so that the Yukawa potential in Equation (35) is attractive at large distances and repulsive at small distances. Depending on the strength of the coupling $g_{v}$ and the amount of captured DM $f_{\mathrm{DM}}$, DM can form a core (unstressed points), a halo (shadowed yellow points), or both along an entire M–R curve. A transition between a DM core and halo occurs as $f_{\mathrm{DM}}$ increases.
Increasing the coupling $g_{v}$ and the DM fraction $f_{\mathrm{DM}}$ enhances the amount of DM settling in the NS’s core and the star’s self-attraction intensifies. The same happens with increasing scalar coupling $g_{s}$; similar effects of $g_{s}$ and $g_{v}$ on DANSs are anticipated because the scalar and the vector contributions in Equations (32) and (33) share the same form and combine additively. Since it is assumed that DM does not interact with BM, it does not significantly contribute to the degeneracy pressure that counters the contraction. The weak interactions among DM fermions yield a less dense packing compared to BM within NSs. The resulting degeneracy pressure adds a negligible contribution to the BM. Consequently, the outward pressure can support a lower mass, leading to DANSs that are less massive and more compact (i.e., with smaller radii). Note that the radius $R_{\mathrm{DANS}}$ of the DANS corresponds to the outermost radius. If a DM core forms, then $R_{\mathrm{DANS}}=R_{\mathrm{BM}}$ because the DM remains entirely within the star. Otherwise, if a DM halo surrounds the star, $R_{\mathrm{DANS}}=R_{\mathrm{DM}}$ because DM extends beyond the “baryonic region”. In both scenarios, however, the visible radius still remains $R_{\mathrm{BM}}$.
Figure 4: Left: Mass–radius relations for different equations of state of BM (solid lines). The gray-shaded regions are excluded due to GR ($R>2GM/c^{2}$), finite pressure ($R>9GM/(4c^{2})$), and causality ($R>8GM/(3c^{2})$) bounds Bucciantini , respectively. The rotation of the fastest spinning pulsar, J1748-2446ad Hessels et al. (2006), excludes the yellow-shaded area, whereas the red and orange strips correspond to the masses of the two heaviest pulsars: J0740+6620 Cromartie et al. (2019); Fonseca et al. (2021) and J1614-2230 Demorest et al. (2010). The figure is taken from Pastor-Marazuela et al. (2023). Right: Mass–radius diagram when DM accumulates within an NS. The plot is adapted from Grippa et al. (2024), where fermionic, interacting DM described by the Lagrangian in Equation (34) is considered. The black curve corresponds to the BSk22 nuclear EoS, namely, without DM. The other solid lines represent the mass–radius relations with different vector coupling $g_{v}$ and/or DM fractions $f_{\mathrm{DM}}$. Configurations highlighted in yellow represent DANSs with a DM halo, otherwise a DM core forms. The blue strip is drawn for GW190814 Abbott et al. (2020); the orange and gray areas correspond to the estimates of the masses of the NSs involved in GW170817 Abbott et al. (2017b, 2018); the magenta and purple regions come from the 2019 NICER data of PSR J0030+0451 from Miller et al. Miller et al. (2019) and Riley et al. Riley et al. (2019); and the dark and light blue areas represent the 2021 NICER data of PSR J0740+6620 from the same groups Miller et al. (2021); Riley et al. (2021).
According to the EoS in Equations (36) and (37), the interaction with the dark vector field contributes additively to the pressure. As such, if a DM core is formed, the maximum mass slightly increases as $g_{v}$ becomes greater, though always being below the TOV limit in the absence of DM (black curve). This happens because now more pressure can counteract the gravitational force pushing inward. However, if $g_{v}$ and/or $f_{\mathrm{DM}}$ are increased further, DM no longer settle in the core; rather, a halo surrounding the DANS forms. Interestingly, as for bosonic DM Karkevandi et al. (2022), a core–halo transition may occur along the same M–R curve. For given $m_{\Psi}$, $g_{i}$, and $f_{\mathrm{DM}}$, the outermost radius may interchange between $R_{\mathrm{BM}}$ and $R_{\mathrm{DM}}$. Configurations exhibiting a DM halo possess much larger radii, although the maximum mass is compatible. While DM halos do not directly impact the optical radius measured by NICER, their inclusion alters the overall mass and radius profile of the NS, which could affect the compatibility of the model with multi-messenger observational constraints.
Interestingly, a small variation in the coupling is sufficient to produce significant changes in the M–R relations. This feature aligns with findings in Collier et al. (2022). The vector (and scalar) coupling constants are characterized by a narrow interval within which they produce relevant effects. If the couplings are too large, the pressure might become negative, making the EoS nonphysical. Conversely, if the couplings are too small, the interactions would become ineffective, virtually resulting in a free Fermi gas, described by Equations (32) and (33). Moreover, the monotonic increase in the stellar mass with greater coupling(s) emerges because both $\epsilon$ and $P$ are functions of the ratio $c_{i}$.
While the masses of the DM candidates have been considered fixed so far, the effects of a variable $m_{\chi}$ have been demonstrated by Routaray et al. Routaray et al. (2023). Their findings indicate that lighter DM particles ($\sim$300300300300 MeV) are more likely to form extended DM halos, whereas heavier DM candidates tend to be confined to the core of DANSs. Furthermore, they stressed that with intermediate DM mass ($\sim$500500500500 MeV) cores form only with high $f_{\mathrm{DM}}$.
[…]
Such a trend is confirmed even when considering different EoSs for BM. For instance, Ref. Liu et al. (2023) derives the BM EoS from a Brueckner–Hartree–Fock approach, covering all $f_{\mathrm{DM}}\in[0,1]$. Their findings confirm that massive DM particles, low fractions, and weak couplings favor core formation and, in turn, result in smaller and denser DANSs. Additionally, they found a couple of interesting outcomes. First, stable high-$f_{\mathrm{DM}}$ stars can be produced under specific conditions: the hosting NSs must be very light (with mass $\leq 0.5$ $M_{\mathrm{\odot}}\,$if $f_{\mathrm{DM}}\sim 0.9$—see Figure 3 in Liu et al. (2023)) and DM should accumulate through exotic mechanisms such as accretion of normal matter onto a dark star Ciarcelluti and Sandin (2011); Ellis et al. (2018); Gleason et al. (2022). Second, twin DANSs may exist, namely, a low-$f_{\mathrm{DM}}$ star with a DM core and a high-$f_{\mathrm{DM}}$ star with a DM halo both with $M_{\mathrm{DANS}}\sim 0.5$ $M_{\mathrm{\odot}}\,$and $R_{\mathrm{DANS}}\sim 12$ km. These twins would be indistinguishable based solely on mass and optical radius measurements, yet their internal structure and related properties would be fundamentally different (due to the way DM is distributed).
Similar trends are observed in Nelson et al. (2019); Karkevandi et al. (2022); Rafiei Karkevandi et al. (2021) and other studies. A few differences may include the following: the “threshold” DM mass above which only DM cores form (with a realistic radius) and the possible existence of intermediate regions where the behavior of the maximum mass is not constant due to the core–halo transition. For instance, Karkevandi et al. Karkevandi et al. (2022) found that, within their model, if $105\,\mathrm{MeV}<m_{\chi}<200\,\mathrm{MeV}$, then the maximum mass decreases for low fractions because a core is formed, but increases for larger $f_{\mathrm{DM}}$ as a halo forms, potentially exceeding 2 $M_{\mathrm{\odot}}\,$. Conversely, if $m_{\chi}\geq 200$ MeV, only a core forms and, even with very large fractions, the maximum mass never reaches 2 $M_{\mathrm{\odot}}\,$.
We emphasize that for both fermionic and bosonic DM the potential shift in the M–R curve induced by the accumulation of DM could significantly aid in detecting its presence within NSs. In this regard, Ciarcelluti et al. Ciarcelluti and Sandin (2011) demonstrated that the stellar properties of the sources 4U 1608-52 Guver et al. (2010a), 4U 1820-30 Guver et al. (2010b), and EXO 1745-248 Ozel et al. (2010) are more effectively explained by incorporating a heavy DM core (at least 25%) rather than relying solely on pure nucleonic EoSs. However, without additional supporting evidence, the characteristic M–R curves alone cannot be considered definitive proof for the existence of DANSs. Several analyses have found that, within certain ranges of DM parameters, DM models can be degenerate with pure nuclear EoSs. This could occur for several reasons. First, the DM fraction may be so small that no detectable effects arise. Second, two-fluid DM models and nuclear EoSs can both be compatible with observations. Scordino et al. Scordino and Bombaci (2024) estimated that for a typical $1.4M_{\odot}$ NS with a DM fraction of $f_{\mathrm{DM}}=6\%$, the observable stellar radius is reduced by approximately 0.6 km (around 5%). These changes fall within the current observational uncertainties of instruments such as NICER, which measures NSs radii. Also in the model presented here (Figure 4), the black curve, representing a pure NS described by the BSk22 EoS, and all the 5% DM models with $g_{v}\sim 10^{-14}$, are simultaneously consistent with experimental constraints posed from GW170817, PSR J0740+6620, and PSR J0030+0451. Third, the effects of DM on observable properties are not easy to isolate because they may mimic those caused by exotic baryonic states or variations in the nuclear EoS. As described, the softer EoS caused by DM accumulation leads to more compact DANSs with smaller radii, but the same effect might also arise from the inclusion of hyperons Zdunik et al. (2004); Weissenborn et al. (2012); Vidaña (2022) or deconfined quarks Schertler et al. (2000); Baym et al. (2018) in pure nuclear models.
On the other hand, these analyses are not only valuable for strengthening the case for the presence of DM but also for ruling out some combinations of baryonic and DM models. For instance, Ellis et al. Ellis et al. (2018) demonstrated that well-established nuclear EoSs, such as ALF2, APR4, ENG, H4, and SLy4, are incompatible with the mass measurement of PSR J0348+0432 if the accreted DM fraction exceeds 5%.
Figure 5: $\Lambda$–M relations for a DANS with fermionic DM described by the Lagrangian in Equation (34). As in Figure 4, the black curve is obtained when no DM accumulates, whereas changing $g_{v}$ and/or $f_{\mathrm{DM}}$ provides significant modifications. The gray region represents the 90% confidence upper bound from GW170817 Abbott et al. (2019), and the light and dark blue regions correspond to the same bounds for the primary and secondary compact objects of GW190425 Abbott et al. (2020); Yang et al. (2022) (assuming low-spin priors). Plot is adapted from Grippa et al. (2024).
The (dimensionless) tidal deformability $\Lambda$ is computed by means of Equation (23), where $M\equiv M_{\mathrm{DANS}}$ and $R\equiv R_{\mathrm{DANS}}$, i.e., the outermost radius. As for the mass–radius relations, in Figure 5 we report how a nuclear EoS (as BSk22) is modified if DM accumulates within the star according to the afore-mentioned model of fermionic, interacting DM. For a given total gravitational mass, the presence of DM condensed in a core results in a smaller tidal deformability Ivanytskyi et al. (2020); Liu et al. (2024) compared to a pure NS. When a DM halo forms, both the mass and radius of the NS decrease. However, in the scaling relation $\Lambda\propto(M_{\mathrm{DANS}}/R_{\mathrm{DANS}})^{-5}$ (Equation (23)), the reduction in radius has a dominant effect over the reduction in mass, resulting in an overall decrease in tidal deformability. From the perspective of a distant observer, these changes would be interpreted as an effective softening of the EoS. Indeed, with a softer EoS less pressure can be exerted to balance the gravitational self-interaction, so the star becomes more compact with a smaller radius. This increased compactness results in the star being more resistant to deformation, thus lowering its $\Lambda$. However, as pointed out in Giangrandi et al. (2023), this effect may be not easy to identify. As for the M–R curves, the modifications in the $\Lambda$–M relations alone are not sufficient to establish the presence of DM within NSs. Indeed, both the DM capture and the presence of an additional degree of freedom (such as hyperons) can contribute to the softening of the EoS. As a consequence, distinguishing these scenarios based on the measurement of $\Lambda$ and/or the observed radius mass would be challenging. In other words, despite completely different internal structures, the same mass and radius could correspond to two distinct situations: a DANS with a certain fraction of DM and a pure NS containing hyperons, SQM, or other new particles. In contrast, the existence of a DM halo leads to a substantial rise in the tidal deformability because of the significant increase in $R_{\mathrm{DANS}}$.
As for the dependence on the coupling, $\Lambda$ increases monotonically with larger $g_{v}$, eventually favoring the formation of a DM halo. Das et al. Das et al. (2022) have emphasized this evidence is important in differentiating a single-fluid model from a two-fluid approach when describing DANSs. Indeed, from the analysis in Das et al. (2019) (where DM is coupled to BM through a Higgs portal interaction, forming one fluid), the stellar properties, including $\Lambda$, always decrease in the presence of DM. In contrast, in the two-fluid scenario, $\Lambda$ can rise beyond the observational limits when a halo extends over $\sim$20202020 km. Due to such a great radius, the resulting curves do not usually match experimental observations, making configurations with a halo disfavored by these DM models.
3.5 Speed of Sound in Dark Matter Admixed Neutron Stars
The stiffness of matter in compact objects is described by the sound speed $c_{s}$, which plays a critical role in preventing a static NS from collapsing into a BH. Unlike the mass or the radius, the sound speed cannot be directly measured by experiments, nor it is easily inferred. However, the tidal deformability, $c_{s}$, is still a key parameter for characterizing the NS properties and, in particular, for quantifying how resistant the matter is to compression in extreme conditions. In this regard, Ecker et al. Ecker and Rezzolla (2022) have demonstrated that in light stars (whose $c_{s}$ increases gradually towards the center), the outer layers are soft while the core is rather stiff; whereas heavy NSs have a stiffer outer shell and a softer core (where the sound speed reaches a local minimum).
By definition, the sound speed represents the velocity at which linear perturbations propagate in a uniform fluid Rezzolla and Zanotti (2013) and can be derived from the EoS as $c_{s}^{2}=\left(\frac{\partial P}{\partial\epsilon}\right)_{s}$, that is, assuming thermodynamic equilibrium and computing the derivative at constant specific entropy $s$. This definition implies that a stiffer EoS leads to a larger $c_{s}$, whereas smaller sound speeds are obtained with softer EoSs. Since NSs are not uniform spheres of constant density, $c_{s}$ varies as one moves from the surface to the core. Due to the uncertainty about the nuclear EoS governing the interiors of NSs, the exact trend in $c_{s}$ remains a crucial open question for understanding dense matter properties and GW emission from compact objects.
Minimal physical constraints on the sound speed are posed by causality Bludman and Ruderman (1968) and thermodynamic stability, requiring $0\leq c_{s}\leq 1$. Beyond that, theoretical constraints can be derived. At low densities $n\lesssim n_{0}=0.16\,\mathrm{fm^{-3}}$ ($n_{0}$ is the nuclear saturation number density), the condition $c_{s}<<c$ has been well established Hebeler et al. (2013); Altiparmak et al. (2022). Conversely, at high densities ($n>>n_{0}$) exotic states of matter (like free quarks and gluons) may exist, making it hard to predict the profile of $c_{s}$. In this regime, an important open question is whether the so-called sub-conformal limit $c_{s}<c/\sqrt{3}$ should hold for the NS’s EoS Bedaque and Steiner (2015); Hoyos et al. (2016); Landry et al. (2020); Roy and Suyama (2024). Such a limit reflects the restoration of the conformal symmetry of quantum chromodynamics (QCD) at asymptotically large densities Kurkela et al. (2010); Fraga et al. (2014). Hence, in conformal theories, the sound speed is expected to approach the value $c/\sqrt{3}$, realized in ultrarelativistic fluids from below Tews et al. (2018); Greif et al. (2019). Some studies argue for violations of this limit Kojo et al. (2015); Alsing et al. (2018); Bitaghsir Fadafan et al. (2019); McLerran and Reddy (2019), which is further supported by its tension with measurements of the most massive NSs Demorest et al. (2010); Antoniadis et al. (2013); Fonseca et al. (2021) as well as theoretical predictions on the maximum (gravitational) mass Bauswein et al. (2017); Margalit and Metzger (2017); Rezzolla et al. (2018); Ruiz et al. (2018); Shibata et al. (2017, 2019); Nathanail et al. (2021) and from QCD Altiparmak et al. (2022) (and references therein).
DM accumulation inside NSs alters the total pressure and energy density, thus affecting $c_{s}$. While the literature lacks comprehensive analyses of DM’s influence on DANSs’ speed of sound, an important contribution comes from Giangrandi et al. Giangrandi et al. (2023), who derive an effective sound speed for a two-fluid system of (bosonic) DM and BM:
$c_{s\mathrm{,\,eff}}^{2}=\eta c_{s\mathrm{,\,BM}}^{2}+(1-\eta)c_{s\mathrm{,\,DM}}^{2}\,,$ (38)
where $c_{s\mathrm{,\,BM}}^{2}$ and $c_{s\mathrm{,\,DM}}^{2}$ are the speed of sound for the baryonic and the DM fluid, respectively. In Equation (38), $\eta$ is a parameter related to the thermodynamic properties of the two components:
$\eta=\frac{\partial\epsilon_{\mathrm{BM}}}{\partial\mu_{\mathrm{BM}}}\left(\frac{\partial n_{\mathrm{BM}}}{\partial\mu_{\mathrm{BM}}}+\xi^{2}\frac{\partial n_{\mathrm{DM}}}{\partial\mu_{\mathrm{DM}}}\right)^{-1}\,,$ (39)
where $\xi=\mu_{\mathrm{DM}}/\mu_{\mathrm{BM}}$ is the ratio of chemical potentials. This formulation ensures that $\eta\in[0,1]$ and, in turn, $c_{s\mathrm{,\,eff}}^{2}$, lies between the ratios of the pure components. In particular, its lower value is obtained for a pure DM fluid ($\eta=0$); whereas the upper edge, reached with $\eta=1$, corresponds to an “ordinary” NS of only BM.
Figure 6 Giangrandi et al. (2023) clarifies the influence of DM on the NS’s effective sound speed. In detail, if a DM core forms (left panel in Figure 6), then $c_{s\mathrm{,\,eff}}^{2}$ is significantly smaller than $c_{s\mathrm{,\,BM}}^{2}$ and the DANS’s EoS becomes softer overall. As expected, the difference is more pronounced moving towards the core, where DM is accumulated. In contrast, if a DM halo forms (right panel in Figure 6), no significant distinctions arise between $c_{s\mathrm{,\,eff}}^{2}$ and $c_{s\mathrm{,\,BM}}^{2}$ as long as the density is large enough. However, near the crust and up to the surface (i.e., at low energy densities), the DM component begins to dominate, consistently having a halo configuration.
Figure 6: The (squared) effective sound speed $c_{s\mathrm{,\,eff}}^{2}$ versus total energy density $\epsilon_{\mathrm{tot}}$ for DANSs with a DM core (left) or a DM halo (right) for two different values of the DM candidate mass $m_{\chi}$ and the fictional mass $m_{I}$ defined in Equations (30) and (31) according to Giangrandi’s model Giangrandi et al. (2023) (from which the plots are taken). In the scenario of DM core formation, $c_{s\mathrm{,\,eff}}^{2}$ is in between the speed of sound of a pure fluid of BM (red curve, derived from the nuclear IST EoS) and DM (black curve). The effective sound speed decreases with a larger DM fraction (represented by the parameter $\xi$ defined in Equation (39)), and the discrepancy increases with greater energy density, meaning that near the core, where DM is accumulated, the DANS’s EoS is significantly softer. In contrast, the presence of a DM halo makes $c_{s\mathrm{,\,eff}}^{2}$ close to the sound speed of a BM fluid at large densities, while in the crust (i.e., near the surface) the DM begins to dominate, as expected.
A similar analysis by Thakur et al. Thakur et al. (2024) examined fermionic DM accreted through neutron decay (see Section 5). Their findings, along with the previous discussion, show that the presence of DM softens the EoS at intermediate densities, reducing the sound speed compared to a pure BM NS and meaning that DM tends to settle close to the NS’s core.
Before ending this section, we emphasize that significant knowledge about extremely dense matter in compact objects can be derived from the sound speed because its own definition is directly related to the EoS. Analyzing the trend in $c_{s}$ allows for comparison of theoretical models that incorporate DM and those that exclude it with observations in order to constrain the DM parameter space and to support/rule out some EoS frameworks.
However, as for the M–R and $\Lambda$–M relations discussed in Section 3.4, some degeneracy between purely baryonic stars and DANSs may arise. Again, this would manifest because baryonic models, with suitable parameter adjustments, can replicate variations in $c_{s}$ that resemble those predicted in DANSs. Hence, a comprehensive approach that synthesizes both observational (a robust dataset containing information on mass, radius, $\Lambda$, $c_{s}$ and thermal evolution) and theoretical insights is needed to resolve this ambiguity and identify the distinct signatures of DM in NSs.
4 Dark Matter Signal in Gravitational Wave Emission
As DM accumulates in NSs, it can alter their structure and potentially leave a distinct signature in the GW emission from binary systems observed by ground-based detectors such as Virgo Acernese et al. (2015), LIGO Aasi et al. (2015), and KAGRA Somiya (2012). Analytical studies Ellis et al. (2018); Kopp et al. (2018); Bhattacharya et al. (2023); Croon et al. (2018); Choi and Jung (2019); Alexander et al. (2018) have computed the contribution of DM to the GW spectrum. If DANSs involved in coalescence contain significant amounts of DM, interaction between their dark components via long-range forces may occur. To remain as model-independent as possible, these interactions can be parametrized by a Yukawa-type potential Alexander et al. (2018), expressed as $V_{Y}(r)=\alpha\frac{m^{2}\eta}{r}e^{-r/l}$, where $l=m_{\chi}^{-1}$ is the length scale of the interaction, $m=m_{1}+m_{2}$ is the binary’s total mass, $\eta=m_{1}m_{2}/m$ the reduced mass of the inspiraling compact objects, $r$ the orbital separation, and $\alpha=\pm(q_{1}q_{2})/(m_{1}m_{2})$ represents the strength of the potential relative to gravity (with the positive sign in case of an attractive force, whereas the negative sign means repulsion). During the (early) inspiral, the DANSs can be modeled as point masses/charges and, as long as the distance between the companions greatly exceeds $l$, the system’s GW emission follows the predictions of GR. However, once their relative distance dropped below the range of that force, the Yukawa interaction would manifest and modify the potential energy of the binary system. This yields a couple of important observable effects. First, detectors would observe a modification in the chirp mass $\mathcal{M}=\eta^{3/5}(m_{1}+m_{2})^{2/5}$. The additional dark force influences the rate of orbital decay by affecting the time evolution of the system’s frequency and the time of the merger. This, in turn, would cause deviations in the GW waveforms that are degenerate with a shift in the DANS’s masses. This degeneracy is further explored in Croon et al. (2018), where it is demonstrated that dark photons with masses $\mathcal{O}(10)\;\mathrm{km}\lesssim(m_{\chi})^{-1}\lesssim\mathcal{O}(1000)\;\mathrm{km}$ emitted by inspiraling NSs would manifest as a time-dependent evolution of the chirp mass, quantified as $\sim$$(1-\alpha^{\prime})^{2/5}$, with $\alpha^{\prime}\sim 1/3$ Sagunski et al. (2018) lying within the LIGO-Virgo sensitivity band. Second, a dipole radiation may emerge and influence the period evolution of companion stars with significantly asymmetric charge-to-mass ratio. This occurs once the orbital angular frequency of the binary exceeds $l^{-1}$, affecting the inspiral Alexander et al. (2018); Hook and Huang (2018); Kopp et al. (2018); Poddar et al. (2024).
For such effects to be detectable and distinguishable from ordinary NS mergers, significant DM accumulation is required. This cannot be achieved through gravitational capture alone (Section 2), which typically results in DM cores no larger than $\leq 10^{-10}M_{\odot}$. Indeed, assuming a DM candidate with $m_{\chi}\lesssim 1$ GeV and ambient density $\rho_{\chi}\sim 1\;\mathrm{GeV/cm^{3}}$, the total number of DM particles (computed through Equation  (2.1) or Equation (6) in McDermott et al. (2012)) accreted solely via gravitational capture is much smaller than the baryon number $N_{\mathrm{B}}$ in NSs, i.e., $N_{\chi}\sim 10^{47}<<10^{57}\sim N_{\mathrm{B}}$. To obtain a much higher value that is consistent with the DM fractions assumed in Section 3.4, either DM is accreted in much denser regions or alternative accumulation mechanisms must be taken into account.
By assuming that any DM particle that passes through an NS over its lifetime is captured, Kopp et al. Kopp et al. (2018) established upper limits on the number $N_{\chi}^{\mathrm{capt}}$ of DM particles accreted in an NS and on the coupling $\alpha\lesssim 10^{-15}$ that leads to an almost negligible Yukawa force. To obtain much larger values of $N_{\chi}^{\mathrm{capt}}$ and $\alpha$ and, in turn, to observe meaningful effects, the assumption that DM can be accreted only through direct capture should be relaxed, and more exotic accumulation mechanisms need to be considered. One proposal involves DM production through bremsstrahlung in a newly formed hot NS Nelson et al. (2019); Ellis et al. (2018). Another possibility is the generation of DM via hypothetical decay of SM particles. Section 5 is dedicated to this topic.
Additional phenomena have been investigated in scenarios where DANS merge. Ellis et al. Ellis et al. (2018) found that the power spectral density of emitted GWs could exhibit additional peaks beyond the primary mode $f_{2}$, typically seen in the range 2–4 kHz Bauswein and Janka (2012); Clark et al. (2016); Prakash et al. (2021, 2024); though similar extra peaks have sometimes been found also in NS merger simulations, where secondary oscillation modes can emerge due to non-linear interaction between the fundamental quadrupolar mode $f_{2}$ and quasi-radial oscillations of the remnant Stergioulas et al. (2011); Takami et al. (2014); Bauswein et al. (2016). Hook et al. Hook and Huang (2018) highlighted that a force mediated by light DM particles, such as axions, could impact the inspiral phase and be detected. Bezares et al. Bezares et al. (2019) observed a couple of interesting characteristics: In DANS mergers, the three-body interaction between the two DM cores and the fermionic matter may trigger a pronounced one-arm instability of the remnant, thus breaking the traditional quadrupolar symmetry and forming an $|m|=1$ over-density. The gravitational waveform would then exhibit an additional $|m|=1$ mode, occurring at half the frequency of the dominant mode $|m|=2$, which, in turn, would decay more rapidly over time compared to standard BNS coalescences. Remarkably, while these latter can sometimes mimic similar features (especially in asymmetric binaries with high spin and/or eccentricity Paschalidis et al. (2015); East et al. (2016); Lehner et al. (2016a, b); Radice et al. (2016)), some features might break this degeneracy. In DANS mergers, the one-arm instability develops even in equal-mass, non-spinning systems, a behavior almost absent in standard NS binaries. Both kinds of coalescence may show the same exponential decay of the dominant $|m|=2$ mode, but when DM is accreted its attenuation is fast enough to be detectable.
Mergers between NSs and their mirror counterparts (containing mirror DM Hodges (1993); Berezhiani et al. (2001); Berezhiani (2004); Ignatiev and Volkas (2003); Foot (2014)) have been analyzed in Hippert et al. (2023). The fate of the remnant depends on $f/v$, namely, the ratio of the vacuum expectation values of standard matter and mirror-sector Higgs fields, respectively. For $f/v\geq 5$, a stable (mirror) DANS may form from the coalescence; whereas smaller ratios yield prompt or retarded collapse of the remnant to a BH. In general, a numerical relativity approach is needed to establish the features of such a merger. Numerical simulations of DANS mergers have been successfully carried out in Bezares et al. (2019); Emma et al. (2022); Bauswein et al. (2023), along with the construction of quasi-equilibrium configurations Rüter et al. (2023) (where it is also shown that bosonic DM halos can induce a “mass-shedding” effect in binary systems). Furthermore, dark photons Diamond and Marques-Tavares (2022) and axions Diamond et al. (2024) produced in these collisions might convert to observable $\gamma$-rays.
Overall, these advancements open a new avenue for studying the dynamics of DANS coalescence and obtaining waveforms to compare with observations. By incorporating DM into models, researchers are now equipped to investigate its potential signatures in GW signals and to test DM theories with current interferometers like Advanced Virgo Acernese et al. (2015), Advanced LIGO Aasi et al. (2015), and KAGRA Somiya (2012) and the next-generation detectors such as the Einstein Telescope Hild et al. (2011); Branchesi et al. (2023), the Cosmic Explorer Abbott et al. (2017); Evans et al. (2023), and LISA Amaro-Seoane et al. (2017); Arun et al. (2022); Seoane et al. (2023). Future detectors may also reveal the influence of extended DM halos on NSs’ tidal deformability and second Love number Dengler et al. (2022).
5 Neutron Decay to Dark Matter
DM interactions can be constrained by investigating neutron decay within NSs. The longstanding discrepancy in neutron decay lifetime, observed using the beam and bottle methods Berezhiani and Bento (2006); Serebrov et al. (2008); Fornal and Grinstein (2018), provides an avenue to explore these interactions. In the bottle method, neutrons are trapped, making it sensitive to the total neutron width, yielding a slightly shorter lifetime of $\tau_{n}^{\mathrm{bottle}}=879.6\pm 0.6~{}\mathrm{s}$ Pichlmaier et al. (2010); Serebrov et al. (2005); Steyerl et al. (2012); Arzumanov et al. (2015). In contrast, the beam method measures the number of protons produced through $\beta$-decay in a fixed volume from a neutron beam, resulting in a longer lifetime of $\tau_{n}^{\mathrm{beam}}=888.0\pm 2.0~{}\mathrm{s}$ Byrne et al. (1990); Yue et al. (2013). This discrepancy can be reconciled by introducing a new decay channel for the neutron into a light dark baryon $(\chi)$, as $n\rightarrow\chi+\cdots$, where the additional particles ()($\cdots$)( ⋯ ) conserve energy and momentum. Consequently, the branching fraction for this decay mode becomes
$f_{n\rightarrow\chi}=1-\frac{\tau^{\mathrm{bottle}}_{n}}{\tau^{\mathrm{beam}}_{n}}=(0.9\pm 0.2)\times 10^{-2}\,.$ (40)
However, more precise measurements of the neutron axial-vector coupling constant $g_{A}$ place stringent limits on the non-standard branching fraction of neutron decay, restricting it to less than $2.7\times 10^{-3}$ at the $95\%$ C.L Czarnecki et al. (2018).
[…]
6 Conclusions
In this review article, we have described the role of DM in NSs, focusing on how its accumulation can influence the structure, evolution, and observable properties of these compact objects. NSs, with their extreme densities and gravitational fields, provide unique environments where DM could accumulate over time. The mechanisms of DM capture and accretion reveal that these compact objects can gather significant amounts of DM, which could lead to heating effects due to self-annihilating interactions. The cooling profile of NSs may be significantly affected by DM scattering and/or annihilation processes. In the latest theoretical framework that incorporates internal mechanisms such as rotochemical heating, the intrinsic heating of NSs turns out to be degenerate with the DM contribution. This degeneracy may potentially be resolved in NSs with long periods Hamaguchi et al. (2019) once more sophisticated methods to evaluate the nucleon pairing and to determine the initial period are developed. Furthermore, if enough DM settles in an NS’s core, it could even trigger the formation of a BH, which in turn could evaporate due to Hawking radiation or consume the whole NS in a relatively short time.
Modeling DANSs often needs a two-fluid framework, where the baryonic and the dark sectors are treated as separate fluids, interacting only through gravity. This approach allows the exploration of different DM candidates, including bosonic and fermionic DM, and the verification of how the formation of DM cores or halos alters observable properties like mass–radius relations and tidal deformability. These properties are closely tied to observable signatures, particularly by GWs produced in BNS mergers. Hence, these models are suitable to be effectively constrained by current interferometers such as LIGO, Virgo, and KAGRA, as well as future third-generation detectors like the Einstein Telescope and Cosmic Explorer. A brief summary is as follows:
  • The formation of a DM core yields a softer EoS and results in a more compact DANS, whose corresponding mass and radius are smaller than those of a star made entirely of BM. In contrast, if DM forms a halo, the EoS turns out to be stiffer, resulting in a larger radius and tidal deformability and slightly greater gravitational mass. In the latter scenario, since $R_{\mathrm{DANS}}$ represents the outermost radius, then $R_{\mathrm{DANS}}=R_{\mathrm{DM}}$ and it can extend significantly beyond the BM. As a consequence, configurations with extended halos are unlikely to match observational constraints and the corresponding EoSs are usually disfavored.
  • Weak interaction couplings, low fractions of DM, and large masses of DM candidates favor the formation of a core.
  • Despite the degeneracy among DM and pure nuclear models, the analyses of M–R and $\Lambda$–M relations can help to rule out certain combinations of DM and baryonic EoSs, while also possibly supporting the evidence for the existence of DANSs.
  • The tidal deformability $\Lambda$ (computed through Equation (23)) is typically reduced in the presence of a DM core, as the star becomes more compact and it is harder for it to be deformed. In contrast, if a DM halo forms, $\Lambda$ increases significantly due to the larger effective radius. This is particularly relevant for GW signals from NS mergers, as the tidal interactions between merging stars leave a distinct imprint on the waveforms.
DANSs may possibly be distinguished in merging binaries. The exertion of a dark force between the dark components of the stars would manifest both in a shift of the masses of the stars that would alter the chirp mass of the system and in a dipole radiation that would reduce the binding energy and influence the binary period evolution.
A fundamental quantity characterizing NSs is $c_{s}$. While it has been widely analyzed for baryonic NSs, the literature lacks a similar comprehensive study of DM’s effect on the sound speed in a DANS. However, an effective sound speed has been derived in Giangrandi et al. (2023). We review how DM cores yield a softer EoS and smaller $c_{s}$, whereas the distribution of DM in a halo favors a larger speed of sound.
An alternative/additional mechanism of DM accumulation in NSs could be the decay of SM particles into DM. The observed $\sim$8888 s discrepancy between beam and bottle neutron lifetime measurements suggests a possible decay channel involving a dark particle $\chi$ with a branching fraction constrained to be $f_{n\rightarrow\chi}\lesssim 2.7\times 10^{-3}$ (90% CL). In particular, a fermionic DM candidate with mass $m_{\chi}\gtrsim 1.2\,\mathrm{GeV}$ or exhibiting strong self-repulsion could resolve that tension.
In conclusion, our review of DANSs has shown that the inclusion of DM in NS models significantly expands the range of possible configurations and observable signatures. As future GW detectors improve and more data become available, matching theoretical predictions with these observations will be critical for confirming the role of DM in compact objects.
[…]
F. Grippa, T. K. Poddar, G. Lambiase conceived the idea of the project. F. Grippa wrote the first version of the manuscript with input from all authors; T. K. Poddar and G. Lambiase supervised the study. All authors have read and agreed to the published version of this review article.
[…]
This research received no external funding.
[…]
No new data were created.
Acknowledgements.
[…]
NS Neutron star
DANS Dark matter admixed neutron star
WD White dwarf
BH Black hole
EoS Equation of state
DM Dark matter
BM Baryonic matter
GW Gravitational wave
BNS Binary neutron star
$\Lambda$ Tidal deformability
SM Standard Model
WIMP Weakly interacting massive particle
CMB Cosmic microwave background
TOV Tolman–Oppenheimer–Volkoff
GR General relativity
M–R Mass–radius
$\Lambda$–M Tidal deformability–mass
BEC Bose–Einstein condensate
RMF Relativistic mean field
SQM Strange quark matter
QCD Quantum chromodynamics
SI Spin independent
SD Spin dependent
CL Confidence level
[…]